OpenStudy (anonymous):

Please help quick! Which of the following is the best estimate of 5 and seven-eighths minus 1 and one-tenth ? Answer 4 and one-half 4 3 and one-half 5

OpenStudy (gw2011):

If you change the fractions to decimals you get: 5 7/8 = 5.875 1 1/10 = 1.1 Now subtract to get: 5.875-1.1 = 4.775 which is approximately 5 Therefore, the best estimate is 5